Three tired and hungry men went to sleep with a bag of apples.
One man awoke, ate (1/3) of the apples, and then went back to
sleep. Later a second man awoke, and ate (1/3) of the
remaining apples, and then went back to sleep. Finally, the third
man awoke and ate (1/3) of the remaining apples. When he
finished there were 8 apples left in the bag. How many apples
were there originally?

Answer by RAY100(1637)

Starting with x total apples,,,each time a man eats he leaves 2/3 of what he found ( eats 1/3,,,1-1/3 = 2/3 )
.
3 men eat therefore 2/3 *2/3 * 2/3 *x =8
,
simplifying, .296296 x = 8, or x =27
.
checking
.
27 *1/3 =9, 27-9 = 18
18 * 1/3 =6,,18-6=12
12*1/3 =4,,,,12-4=8,,,,,ok